OpenStudy (anonymous):

help me i will give u medals Solve the equation. Round the answer to the nearest tenth. 3.2x = 4.6 A. x ≈ 1.4 B. x ≈ 1.3 C. x ≈ 1.5 D. x ≈ 14.7

OpenStudy (anonymous):

plzzz help me

OpenStudy (anonymous):

is it a because to get x you have to divide 3.2 from both side to get x alone, which will give you the answer

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

i have 4 more

OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 1.04x + 3.14 = 7.98 A. x ≈ 4.5 B. x ≈ 10.7 C. x ≈ 4.6 D. x ≈ 4.7

OpenStudy (anonymous):

move 3.14 to the other side so you subtract it from 7.98 which will give you 4.48, and then do the same thing i told from the previous question

OpenStudy (anonymous):

you should get 4.7

OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 5.2x + 3.2 = 9.8 A. 1.1 B. 1.4 C. 1.5 D. 1.2

OpenStudy (anonymous):

1.2

OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 4.12(x – 7.89) = 18.12 A. 12.6 B. 12.5 C. 12.4 D. 12.2

OpenStudy (anonymous):

i got 9.4 as the answer

OpenStudy (anonymous):

9.4 not on there

OpenStudy (anonymous):

is says just estimate so just choose the lowest value

OpenStudy (anonymous):

12.2

OpenStudy (anonymous):

Which estimate is most reasonable? Francesca purchases a box of cereal for $2.98, hamburger for $4.12, tomatoes for $3.07, and milk for $1.96. She has $17 with her. About how much change should she expect to receive? A. $7 B. $6 C. $5 D. $4

OpenStudy (anonymous):

5 dollars

OpenStudy (anonymous):

tyvm

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

u made me fail

OpenStudy (anonymous):

i failed

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

idk i got 60%

OpenStudy (anonymous):

u didnt fail

OpenStudy (anonymous):

yes i did it onley take higher than 80%

OpenStudy (anonymous):

i got below it

OpenStudy (anonymous):

how many question were there?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

i missed 2

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

brbrbrrbrbrbrbrbrbrbrb